# Saner: Annotation-free Societal Attribute Neutralizer for Debiasing CLIP

### v1

@@ 0.B.2 Analysis on the data size
Appendix 0.C List of Person-, Gender-, Age-, Race-Specific Terms
@@
Appendix 0.D List of concepts and occupations
@@
Adjectives: kind, successful, noncriminal, nonviolent, moral, rich, honest, smart, evil, unkind, unsuccessful, criminal, violent, immoral, poor, dishonest, dumb, clever, stupid, hardworking, lazy, nasty, trustworthy, untrustworthy, caring, dangerous, heroic, villainous, good. We refer these concepts from [4]. Regardig the templates used to create prompts, we use “This is the face of a {} person”, “This is a photo of a {} person”, “This person is {}”, “A {} person”, “A {} individual”, “This individual is {}”, “Photo of a {}”, “{}”, “This is a {}”, “This is the photo of a {}”, “A {}”, following [4]. The curly brackets are replaced with each concept.
Occupations: doctor, aerospace engineer, computer programmer, electrical engineer, scientist, artist, author, designer, musician, painter, photographer, singer, writer, architect, civil engineer, engineer, programmer, software developer, childcare worker, coach, dental assistant, dental hygienist, dentist, clerk, housekeeper, maid, massage therapist, nurse, therapist,psychologist, social assistant, social worker, teacher, professor, CEO, skateboarder, surfer, baseball player, football player, soccer player, tennis player. For the occupation list, we refer to [16] with some modifications such as additional occupations for more inclusive list. As for the prompt templates, we select ones that are suitable to occupations (e.g., “Photo of a {}”). We use this list for the text-to-image experiments with some modifications. Specifically, we remove the similar occupations (e.g., remove civil engineer as there is engineer). Additionally, we remove occupations where the text-to-image model [39] can not generate images with humans or it tends to generate multiple individuals.
Activities: arranging flowers, playing tennis, playing skateboarding, playing baseball, playing soccer, playing football, playing snowboarding, playing skiing, cleaning, dressmaking, tying tie, smiling, crying, laughing, cooking, making pizza, dancing, drinking beer, drinking wine, eating hotdog, eating cake, using computer, playing game, gardening, singing, petting dog, petting cat, makeup, shopping, playing piano, playing guitar, carrying baby. For activities, we use a subset of the Kinetics dataset [25] with some additional activities. For the prompt templates, we use “This is the face of a person who likes {}”, “ This is a photo of a person who likes {}”, “This person likes {}”, “A person who likes {}”, “Photo of a person who likes {}”, “This is a person who likes {}”.
Appendix 0.E Additional visual examples for image generation experiments
@@
Appendix 0.F Potential Negative Impact

### v3

@@ B.3 Analysis on the data size
B.4 Quality of the generated images
We evaluate image fidelity and image-text alignment on the COCO Karpathy test set. Specifically, we use FID score (Heusel et al., 2017) and CLIPScore (Hessel et al., 2021) to measure fidelity and image-text alignment, respectively. The results show that SANER achieves comparable performance to the original Stable Diffusion (FID: 28.1, CLIPScore: 31.2 vs. FID: 28.7, CLIPScore: 31.0 for SANER), confirming that SANER achieves debiasing while maintaining image generation quality.
B.5 Experiments on BLIP
To verify the effectiveness of SANER for VLMs beyond CLIP, we conduct gender bias experiments using BLIP (Li et al., 2023a). As shown in Table 9, SANER demonstrates superior debiasing performance compared to the original BLIP and the projection-based debiasing.
Table 9: Gender bias for BLIP, evaluated by MaxSkew@1000.
FairFace PATA
BLIP Model Adjective Occupation Activity Adjective Occupation Activity
Original $16.8$ $15.3$ $12.8$ $7.7$ $7.4$ $5.5$
Projection $\mathbf{14.9}$ $19.8$ $17.4$ $12.0$ $14.7$ $17.4$
SANER (Ours) $15.3$ $\mathbf{13.9}$ $\mathbf{10.9}$ $\mathbf{6.7}$ $\mathbf{6.7}$ $\mathbf{4.9}$
B.6 Experiments on FACET
In addition to FairFace and PATA, we evaluate SANER on the FACET dataset (Gustafson et al., 2023). Similar to PATA, FACET comprises real-world, natural images but includes a broader range of annotations for protected attributes, such as hair color. To evaluate SANER’s effectiveness on FACET, which differs in data distribution from FairFace and PATA, we conducted experiments focusing on gender bias. The results, presented in Table 10, demonstrate SANER’s superior debiasing performance, further supporting its robustness across diverse dataset distributions. Experiments on other attributes, such as hair color, are left for future work.
Table 10: Gender bias on FACET, evaluated by MaxSkew@1000.
FACET
CLIP Model Adjective Occupation Activity
Original $46.0$ $37.7$ $33.4$
Projection $37.3$ $46.3$ $41.1$
SANER (Ours) $\mathbf{31.7}$ $\mathbf{24.7}$ $\mathbf{25.2}$
Appendix C List of Person-, Gender-, Age-, Race-Specific Terms
@@
Appendix D List of concepts and occupations
@@
Adjectives: kind, successful, noncriminal, nonviolent, moral, rich, honest, smart, evil, unkind, unsuccessful, criminal, violent, immoral, poor, dishonest, dumb, clever, stupid, hardworking, lazy, nasty, trustworthy, untrustworthy, caring, dangerous, heroic, villainous, good. We refer these concepts from (Berg et al., 2022). Regardig the templates used to create prompts, we use “This is the face of a {} person”, “This is a photo of a {} person”, “This person is {}”, “A {} person”, “A {} individual”, “This individual is {}”, “Photo of a {}”, “{}”, “This is a {}”, “This is the photo of a {}”, “A {}”, following (Berg et al., 2022). The curly brackets are replaced with each concept.
Occupations: doctor, aerospace engineer, computer programmer, electrical engineer, scientist, artist, author, designer, musician, painter, photographer, singer, writer, architect, civil engineer, engineer, programmer, software developer, childcare worker, coach, dental assistant, dental hygienist, dentist, clerk, housekeeper, maid, massage therapist, nurse, therapist,psychologist, social assistant, social worker, teacher, professor, CEO, skateboarder, surfer, baseball player, football player, soccer player, tennis player. For the occupation list, we refer to (Friedrich et al., 2023) with some modifications, such as additional occupations for a more inclusive list. As for the prompt templates, we select ones that are suitable to occupations (e.g., “Photo of a {}”). We use this list for the text-to-image experiments with some modifications. Specifically, we remove the similar occupations (e.g., remove civil engineer as there is engineer). Additionally, we remove occupations where the text-to-image model (Rombach et al., 2022) can not generate images with humans or it tends to generate multiple individuals.
Activities: arranging flowers, playing tennis, playing skateboarding, playing baseball, playing soccer, playing football, playing snowboarding, playing skiing, cleaning, dressmaking, tying tie, smiling, crying, laughing, cooking, making pizza, dancing, drinking beer, drinking wine, eating hotdog, eating cake, using computer, playing game, gardening, singing, petting dog, petting cat, makeup, shopping, playing piano, playing guitar, carrying baby. For activities, we use a subset of the Kinetics dataset (Kay et al., 2017) with some additional activities. For the prompt templates, we use “This is the face of a person who likes {}”, “ This is a photo of a person who likes {}”, “This person likes {}”, “A person who likes {}”, “Photo of a person who likes {}”, “This is a person who likes {}”.
Appendix E Additional visual examples for image generation experiments
@@
Appendix F Potential Extension
Additional Attributes. While we have aimed to make the attribute list as comprehensive as possible based on prior work (Berg et al., 2022; Seth et al., 2023; Chuang et al., 2023), we acknowledge that the choice of attribute groups may be subjective and that certain attributes might be missed. However, in contrast to previous approaches that require extensive attribute labels for images in the dataset, our method is designed to be flexible and extensible, allowing the attribute list to be expanded or adjusted based on specific needs or ethical considerations of the application domain.
Non-binary gender. While our experiments followed prior work (Berg et al., 2022; Seth et al., 2023; Chuang et al., 2023) in focusing on binary gender, SANER naturally extends to non-binary gender by defining additional attribute-specific terms and their neutral forms (e.g., “non-binary person” $\rightarrow$ “person”). The debiasing loss (Eq. 10) can handle any number of attribute groups, making it straightforward to include non-binary gender categories.
Automated text neutralization. While our current word-level approach is effective for well-defined social biases, it could be extended to handle more complex cases. As a potential future direction, embedding-based neutralization or context-aware language models could be explored to automate this process. These methods would enable bias identification and neutralization at a semantic level, reducing reliance on predefined attribute word lists and making the debiasing process more robust and adaptable to diverse scenarios.
Debiasing LLaVA-like models. SANER can be extended to mitigate societal biases in large vision-language models, such as LLaVA (Liu et al., 2024a), by incorporating a debiasing mechanism for the image encoder. Specifically, this could involve training a debiasing layer to remove attribute information from the visual features, particularly for images without human subjects. We leave this extension as an avenue for future work.
Appendix G Potential Negative Impact